OpenStudy (countrygirl1431):

Please help me Please!!! Write the standard form of the line that contains a slope of 2/3 and passes through the point (1, 1). Include your work in your final answer. Type your answer in the box provided or use the upload option to submit your solution.

OpenStudy (countrygirl1431):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

what do you have so far? were you able to get started?

OpenStudy (countrygirl1431):

no i have no clue what to do

jimthompson5910 (jim_thompson5910):

have you learned about point slope form?

OpenStudy (countrygirl1431):

yes i know how to do point slope form

jimthompson5910 (jim_thompson5910):

ok what is point slope form in general?

OpenStudy (countrygirl1431):



jimthompson5910 (jim_thompson5910):

that's the slope formula

jimthompson5910 (jim_thompson5910):

this is point slope form \[\Large y-y_1 = m(x - x_1)\]

jimthompson5910 (jim_thompson5910):

look familiar?

OpenStudy (countrygirl1431):

yeah

jimthompson5910 (jim_thompson5910):

what is m?

OpenStudy (countrygirl1431):

its the slope you use the other equation that i gave to get m

jimthompson5910 (jim_thompson5910):

well they provided the slope in the instructions

OpenStudy (countrygirl1431):

right i just dont know how to plug the numbers in to the formula the turn it into standard form

jimthompson5910 (jim_thompson5910):

so first we replace m with 2/3 \[\Large y-y_1 = m(x - x_1)\] \[\Large y-y_1 = \frac{2}{3}(x - x_1)\]